### v1

# Symbolic Powers via Extension

Abstract.
This article investigates under which conditions the symbolic powers of the extension of an ideal is the same as the extension of the symbolic powers. Our result generalizes the known scenarios. As an application, we prove formulas for the resurgence of sum of two homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is algebraically closed. Initially, these were known for ideals in polynomial rings.
Key words and phrases:
Resurgence, asymptotic resurgence, prime extension, flat, faithfully flat
2010 Mathematics Subject Classification:
14N20, 13F20, 14C20
1. Introduction
For an ideal $I$ in a noethrian ring $R$, the $m$-th symbolic power is defined as follows:
$I^{(m)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{m}R_{\mathfrak{p}}\cap R.$
The symbolic powers are very well studied objects. The behavior of symbolic powers under homomorphism are also well studied. If $\phi:A\to B$ is a homomorphism of rings, then for any ideal $I$ in $A$ we use $IB$ to denote the extension of the ideal $I$ in $B$ and for any ideal $J$ in $B$ we use $J\cap A$ to denote the contraction of the ideal $J$ in $A$.
In [GHMN17, Proposition 3.8], it has been proved that if $\phi:A={\mathbbm{k}}[y_{1},\dots,y_{s}]\to B={\mathbbm{k}}[x_{0},\dots,x_{n}]$ is defined by $y_{i}\mapsto f_{i}$, with every subset of $c+1$ many $f_{i}$ forms $B$-regular sequence then $I^{(m)}B=\left(IB\right)^{(m)}$, where $I$ is the Stanley-Reisner ideal of a matroid on $[s]=\{1,\dots,s\}$ of dimension $s-1-c$. In [Ake17, Theorem 4], it has been proved that, if $\phi:A={\mathbbm{k}}[y_{0},\dots,y_{n}]\to B={\mathbbm{k}}[y_{0},\dots,y_{n}]$ is given by $y_{i}\to f_{i}$, where $f_{0},\dots,f_{n}$ is an $B$-regular sequence of homogeneous element of the same degree, then $(IB)^{(m)}=(I^{(m)})B$, where $I$ is the defining ideal of a 00-dimensional scheme. In [Wal18, Proposition 2.1], it has been proved that if $\phi:A\to B$ is faithfully flat and the associative primes of $I$ extends to associative primes of $IB$ then the symbolic powers of the extension is the extension of that of $I$.
Motivated by these, in this article, we investigate the following question.
Question 1.1.
Let $A$ and $B$ be noetherian rings, and $I$ be an ideal of $A$. Let $\phi:A\to B$ be a ring homomorphism. When is $(IB)^{(m)}=I^{(m)}B$, where $m\geqslant 2$?
In [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1], the key tool to prove symbolic power extends to symbolic power was that the morphisms were faithfully flat. We proved that one can relax faithfully flatness and prove something more general. We prove the following result extending the known situations where symbolic powers of the extension is the extension of the symbolic powers.
Main Result 1 (Theorem 2.5).
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$, and every associated prime of $I$ is a contraction of some ideal in $B$. Then $(IB)^{(m)}=I^{(m)}B$.
In Remark 2.6, we discussed that our hypothesis is generalizes the known situations. Thus it covers more classes and will broader applications.
One of the central topic of interest in the study of symbolic powers is to find $i$ and $j$ for which $I^{(i)}\subset I^{j}$. In [ELS01, HH02, MS18], it has been proved that if $I$ is a homogeneous ideal of bigheight $h$ in a polynomial ring then $I^{(rh)}\subset I^{r}$ for all $r\geqslant 1$.
In the other hand, one can investigate non-containment between the symbolic and ordinary powers of ideals. The resurgence, defined by Bocci and Harbourne [BH10], and the asymptotic resurgence, defined by Guardo, Harbourne, and Van Tuyl [GHVT13] measure such non-containment.
If $I$ is a homogeneous ideal in a finitely generated ${\mathbbm{k}}$-algebra domain $A$, then the resurgence number, $\rho(I)$ and the asymptotic resurgence, $\rho_{a}(I)$ are defined as follows:
$\rho(I)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subseteq I^{r},m\geqslant 1,r\geqslant 1\right\},$
${\rho_{a}}(I)=\sup\left\{\frac{m}{r}~{}|~{}I^{(mt)}\not\subseteq{I^{rt}},m\geqslant 1,r\geqslant 1,\text{ and }t\gg 1\right\}.$
If $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$ are ideals in polynomial rings ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ then the resurgence and the asymptotic resurgence numbers of $I+J\subset{\mathbbm{k}}[\underline{x},\underline{y}]$ attracted huge attention.
Motivated by 1.1 we study the following question.
Question 1.2.
Let $A$ and $B$ be noetherian rings, $I$ be an ideal of $A$, and $J$ be an ideal of $B$. What can be said about $\rho(I+J)$ and $\rho_{a}(I+J)$, where $I+J=IR+JR$, where $R=A\otimes B$?
The resurgence and the asymptotic resurgence of the sum of ideals $I\subset{\mathbbm{k}}[\underline{x}]$ and $J\subset{\mathbbm{k}}[\underline{y}]$, where ${\mathbbm{k}}[\underline{x}]$ and ${\mathbbm{k}}[\underline{y}]$ are polynomial rings, are studied in [BHJT21, KNT24]. As an application of Theorem 2.5 and Corollary 3.4, we extend those results ideals in standard graded finitely algebra domains over algebraically closed fields.
Main Result 2.
Let $I\subset A$ and $J\subset B$ be homogeneous ideals in finitely generated ${\mathbbm{k}}$-algebra domains $A$ and $B$, where ${\mathbbm{k}}$ is algebraically closed. Set $I+J=IR+JR$, where $R=A\otimes B$. Then the following hold .
  1. (1)
    (Theorem 4.5) $\rho_{a}(I+J)=\overline{\rho}_{a}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
  2. (2)
    (Theorem 4.6) $\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$ Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
The paper is outlined as follows. In Section 2, we prove our main results regarding the relation symbolic powers of an ideal with its extension. In Section 3 we study prime extension property and its applications. In Section 4, we apply results from Section 2 and Section 3 we extend results regarding the sum of ideals in polynomial rings to finitely generated algebra domains over algebraically closed fields.
Acknowledgments.
The authors are thankful to Tài Huy Hà for suggesting the problem and valuable comments. This is a part of the second author’s Ph.D. thesis.
2. Symbolic Powers via flat extension
In this section we prove that with specific conditions, the symbolic powers of an extension is the extension of the symbolic powers. Before proceeding to the main theorem we need to prove the following lemma regarding flat morphisms.
Lemma 2.1.
Let $M$ be a $R$-module where $R$ is a commutative ring, and $I$ be an ideal of $R$, then $(R/I)\otimes_{R}M\cong M/IM$. Furthermore, if $M$ is flat over $A$, we also have $I\otimes_{R}M\cong IM$.
Proof.
The first part can be found in a standard book, e.g., Dummit and Foote. But we provide the proof for completion. After tensoring the exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$ by $M$, we get $I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. It’s a well-known fact that $R\otimes_{R}M\cong M$, then it induces a module homomorphism $\phi$ from $I\otimes M$ to $M$ defined by $\phi:I\otimes M\to M,\phi(r\otimes m)=rm$. Now the image of $\phi$ is $IM$. Since $R\otimes_{R}M\cong M$, the kernel of induced homomorphism from $M$ to $(R/I)\otimes M$ is exactly the image of $\phi$, then by isomorphism theorem, we have $M/IM\cong(R/I)\otimes M$. Since $M$ is a flat, $R$-module we get the following exact sequence: $0\rightarrow I\otimes M\rightarrow R\otimes M\rightarrow(R/I)\otimes M\rightarrow 0$. Following the previous result, we immediately have $I\otimes_{R}M\cong IM$. ∎
Now we are ready to prove one of the key lemmas.
Lemma 2.2.
If a flat ring homomorphism $\phi:A\to B$ extends a prime $\mathfrak{p}$ to a prime $\mathfrak{p}B$, and $\mathfrak{p}$ is a contraction of some ideals in $B$, then $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$ as $A$-algebra.
Proof.
Applying Lemma 2.1 to $B_{\mathfrak{p}B}$, we have $B_{\mathfrak{p}B}\cong((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Our goal is to show $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$.
Define
$\psi_{1}:((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\rightarrow A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\text{ by }\psi_{1}\left(\frac{a\otimes b}{c\otimes d}\right)=\frac{a}{c}\otimes\frac{b}{d}.$
Similarly define
$\psi_{2}:A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\rightarrow((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)\text{ by }\psi_{2}\left(\frac{a}{c}\otimes\frac{b}{d}\right)=\frac{a\otimes b}{c\otimes d}.$
First we need to show both maps are well-defined.
Since $c\otimes d\notin\mathfrak{p}\otimes_{A}B$, then $c\notin\mathfrak{p}$. We also can consider $c\otimes d=1\otimes cd\notin 1\otimes_{A}\mathfrak{p}B$, so $cd\notin\mathfrak{p}B$. We also know that $c\notin\mathfrak{p}$, this forces $d\notin\mathfrak{p}B$ since $\mathfrak{p}B$ is prime. Therefore $\psi_{1}$ is well-defined.
Now we prove that $\psi_{2}$ is also well-defined. Indeed, $c\notin\mathfrak{p}$ and $d\notin\mathfrak{p}B$, then $\phi(c)d\notin\mathfrak{p}B$ because $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. To see this, let $\mathfrak{p}=\mathfrak{q}\cap A$ for an ideal $\mathfrak{q}$ in $B$. From this we observe that $\mathfrak{p}B\cap A=\left(\left(\mathfrak{q}\cap A\right)B\right)\cap A=\mathfrak{q}\cap A={\mathfrak{p}}$. This implies that $\phi(A)\cap\mathfrak{p}B=\phi(\mathfrak{p})$. Note that $\phi(A)\cap\mathfrak{p}B\supset\phi(\mathfrak{p})$. Let $y\in\phi(A)\cap{\mathfrak{p}}B$. Then $y=\phi(x)$ for some $x\in A$. We show that $x\in{\mathfrak{p}}$. Now since $y\in{\mathfrak{p}}B$ then $x\in{\mathfrak{p}}B\cap A={\mathfrak{p}}$.
Notice that $\psi_{1}$ and $\psi_{2}$ are inverses to each other.
Now we show that $\psi_{1}$ is an $A$-algebra homomorphism, then we are done. Verifying the map commutes with the multiplication is straightforward, we only verify the addition here. Let $\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}$ and $\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}$ be elements in $((A\otimes_{A}B)\backslash(\mathfrak{p}B\otimes_{A}B))^{-1}(A\otimes_{A}B)$. Now
$\psi_{1}\left(\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}\right)+\psi_{1}\left(\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}\right)$ $=\frac{a_{1}}{c_{1}}\otimes\frac{b_{1}}{d_{1}}+\frac{a_{2}}{c_{2}}\otimes\frac{b_{2}}{d_{2}}$
$=\frac{a_{1}c_{2}}{c_{1}c_{2}}\otimes\frac{b_{1}d_{2}}{d_{1}d_{2}}+\frac{c_{1}a_{2}}{c_{1}c_{2}}\otimes\frac{d_{1}b_{2}}{d_{1}d_{2}}$
$=\left(\frac{1}{c_{1}c_{2}}\otimes\frac{1}{d_{1}d_{2}}\right)\cdot(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2}).$
Again,
$\psi_{1}\left(\frac{a_{1}\otimes b_{1}}{c_{1}\otimes d_{1}}+\frac{a_{2}\otimes b_{2}}{c_{2}\otimes d_{2}}\right)$ $=\psi_{1}\left(\frac{(a_{1}\otimes b_{1})\cdot(c_{2}\otimes d_{2})}{(c_{1}\otimes d_{1})\cdot(c_{2}\otimes d_{2})}+\frac{(a_{2}\otimes b_{2})\cdot(c_{1}\otimes d_{1})}{(c_{2}\otimes d_{2})\cdot(c_{1}\otimes d_{1})}\right)$
$=\psi_{1}\left(\frac{1}{c_{1}c_{2}\otimes d_{1}d_{2}}\cdot(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2})\right)$
$=\left(\frac{1}{c_{1}c_{2}}\otimes\frac{1}{d_{1}d_{2}}\right)\cdot\left(a_{1}c_{2}\otimes b_{1}d_{2}+c_{1}a_{2}\otimes d_{1}b_{2}\right).$
The following theorem determines the associated primes of an extension of an ideal.
Theorem 2.3.
[Mat89, Theorem 23.2] Let $\phi:A\to B$ be a homomorphism of Noetherian rings, and Let $E$ be an $A$-module and $G$ a $B$-module. If $G$ is flat over $A$, then we have the following
$\operatorname{Ass}_{B}(E\otimes_{A}G)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(E)}\operatorname{Ass}_{B}(G/\mathfrak{p}G).$
Lemma 2.4.
Let $\phi:A\to B$ be a flat homomorphism. If $I\subset A$ is a ideal then we have $\operatorname{Ass}(IB)=\cup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}({\mathfrak{p}}B)$. Furthermore, if $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then we have $\operatorname{Ass}(I^{(n)}B)\subset\operatorname{Ass}(IB)$ for all $n\geqslant 1$.
Proof.
Since $\phi$ is a flat then $A/I\otimes_{A}B\cong B/IB$. Now by Theorem 2.3,
$\operatorname{Ass}(IB)=\operatorname{Ass}_{B}(A/I\otimes_{A}B)=\bigcup_{\mathfrak{p}\in\operatorname{Ass}_{A}(A/I)}\operatorname{Ass}_{B}(B/\mathfrak{p}B)=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\operatorname{Ass}_{B}({\mathfrak{p}}B).$
Note that $\operatorname{Ass}(I^{(n)})\subset\operatorname{Ass}(I)$ for all $n\geqslant 1$. Since $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$ then using Theorem 2.3 we get
$\operatorname{Ass}(I^{(n)}B)=\bigcup_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}\left(I^{(n)}\right)}\operatorname{Ass}({\mathfrak{p}}^{\prime}B)=\left\{{\mathfrak{p}}^{\prime}B~{}|~{}{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I^{(n)})\right\}\subset\operatorname{Ass}(IB),\text{ as }\operatorname{Ass}\left(I^{(n)}\right)\subset\operatorname{Ass}(I).$
The following proves that if a flat morphism extends associative primes to associative primes and each associative prime is a contraction of some ideal then symbolic powers of the extension of an ideal is the extension of the symbolic powers of the ideal.
Theorem 2.5.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. Assume that $I\subset A$ is a proper ideal such that $\operatorname{Ass}_{B}(IB)=\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}$ and every associated prime of $I$ is a contraction of some ideal in $B$ Then $(IB)^{(n)}=I^{(n)}B$, for all $n\geqslant 1$.
Proof.
Using the definition of symbolic powers of an ideal and assumption, we get
$(IB)^{(n)}=\bigcap_{{{\mathfrak{p}}}\in\operatorname{Ass}_{B}(IB)}((IB)^{n}B_{{{\mathfrak{p}}}}\cap B)=\bigcap_{\{{{\mathfrak{p}}}B~{}|~{}{{\mathfrak{p}}}\in\operatorname{Ass}_{A}(I)\}}((IB)^{n}B_{{{\mathfrak{p}}}B}\cap B).$
Now apply Lemma 2.2, we have $B_{\mathfrak{p}B}\cong A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}$. Then by Lemma 2.1, we have the following:
$(IB)^{n}B_{\mathfrak{p}B}\cap B$ $\cong(I^{n}\otimes_{A}B)(A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
$=(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B})\cap(A\otimes_{A}B)$
(*) $\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}\cap B$
Now we prove the last isomorphism. Observe that $B=A\otimes_{A}B\subset A\otimes_{A}B_{\mathfrak{p}B}$. Applying a property from [Mat89, section 7], which states that if $M_{2}$ is flat over $M_{1}$ and $M_{3}$ is flat over $M_{2}$, then $M_{3}$ is flat over $M_{1}$. Here we consider $M_{3}$ as $B_{\mathfrak{p}B}$, $M_{2}$ as $B$, and $M_{1}$ as $A$, then we can see $B_{\mathfrak{p}B}$ is flat over $A$. Now by [Mat89, Theorem 7.4(I)], we consider $I^{n}A_{\mathfrak{p}}$ and $A$ as submodules of $A_{\mathfrak{p}}$ as $A$-modules. Then $\left(I^{n}A_{\mathfrak{p}}\otimes_{A}B_{\mathfrak{p}B}\right)\cap\left(A\otimes_{A}B_{\mathfrak{p}B}\right)=(I^{n}A_{\mathfrak{p}}\cap A)\otimes_{A}B_{\mathfrak{p}B}\cong(I^{n}A_{\mathfrak{p}}\cap A)B_{\mathfrak{p}B}$, then we intersect it with $B$ as $A\otimes_{A}B\cong B$. Recall that $I^{(n)}=\bigcap_{\mathfrak{p}\in\operatorname{Ass}_{A}(I)}(I^{n}A_{\mathfrak{p}}\cap A)$. Therefore we have:
$(IB)^{(n)}=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}(IB)^{n}B_{{\mathfrak{p}}B}\cap B$
[…]
(2.1) $=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{(n)}B_{{\mathfrak{p}}B}\cap B$
(2.2) $=\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}I^{(n)}BB_{{\mathfrak{p}}B}\cap B$
(2.3) $=I^{(n)}B_{W}\cap B,\text{ where }W=B-\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$
(2.4) $=(I^{(n)}B)B_{W}\cap B$
(2.5) $=I^{(n)}B$
First we prove equality 2.1. Since $\phi$ is flat, then using Lemma 2.1, we get
$I^{(n)}B_{{\mathfrak{p}}B}=\left(\cap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{\mathfrak{p}}^{\prime}\cap A)\right)B_{{\mathfrak{p}}B}=\left(\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}(I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\right)\otimes_{A}B_{{\mathfrak{p}}B}$
$=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)\otimes_{A}B_{{\mathfrak{p}}B})=\bigcap_{{\mathfrak{p}}^{\prime}\in\operatorname{Ass}(I)}((I^{n}A_{{\mathfrak{p}}^{\prime}}\cap A)B_{{\mathfrak{p}}B})=(I^{n}A_{\mathfrak{p}}\cap A)B_{{\mathfrak{p}}B}.$
The equality 2.2 is straight forward.
The proof of equality 2.3 is similar to [Gri21, Remark 2.8]. We add the proof for completion. Note that $I^{(n)}B_{W}\cap B=\{r\in B~{}|~{}rs\in I^{(n)}\text{ for some }s\in W\}$ and thus $I^{(n)}B_{W}\cap B\subset\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Now let $r\in\bigcap_{{\mathfrak{p}}\in\operatorname{Ass}(I)}\left(\left(I^{(n)}B_{{\mathfrak{p}}B}\right)\cap B\right)$. Then for each ${\mathfrak{p}}B$ there is $s\not\in{\mathfrak{p}}B$ such that $rs\in I^{(n)}B$. Thus $I^{(n)}B:r$ is not contained in any ${\mathfrak{p}}B\in\operatorname{Ass}(IB)$. Since the associated primes are finite then by Prime avoidance theorem
$I^{(n)}B:r\not\subset\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B.$
Thus there is a $s\in W$ such that $sr\in I^{(n)}B$. Hence we have the equality as sets.
Again the equality 2.4 is straight forward. Lastly, we prove equality 2.5. Now from [AM94, Proposition 3.11(ii)]
$(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s.$
Now by Lemma 2.4, we have $\operatorname{Ass}\left(I^{(n)}B\right)\subset\operatorname{Ass}(IB)$. Thus whenever $s\notin W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$. Thus by [AM94, Proposition 4.4(iii)] we get
$(I^{(n)}B)B_{W}\cap B=\bigcup_{s\in W}I^{(n)}B:s=I^{(n)}B.$
Remark 2.6.
In Theorem 2.5 we extend [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1] by replacing the faithfully flat condition with weaker conditions. Note that faithfully flatness will implies the assumptions in Theorem 2.5. But not vice versa.
We give simple examples satisfying assumptions in Theorem 2.5 but not in [Wal18, proposition 2.1]. Let $A$ be a ring having at least two different prime ideals such that one is not contained in another one, and $B=A_{\mathfrak{p}}$ be localization of $A$ at an arbitrary prime $\mathfrak{p}$. It is a well-known fact that localization is an exact functor, so $B$ is a flat $A$-module, and prime ideal $\mathfrak{p}$ in $A$ is the contraction of $\mathfrak{p}B$. Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [AM94, ex 3.16].
One can take $A={\mathbb{Z}}$ and $B$ as $\mathbb{Z}_{(p)}$ where $p$ is a prime number. Take $I=(p^{m})$ for some $m\geqslant 1$. Then $I$ satisfies assumptions in Theorem 2.5. Furthermore, letting $B$ be any localization of $A$ will meet the assumptions in Theorem 2.5. This tells us that the stalk of $\operatorname{Spec}A$ gives us enough information of symbolic power of an ideal generated by the corresponding point in $\operatorname{Spec}A$.
Another example would be take $A={\mathbb{C}}[t]$ and $B={\mathbb{C}}[t]_{(p(t))}$, where $p(t)$ is an irreducible polynomial in ${\mathbb{C}}[t]$.
3. Resurgence and Prime extension
The following is from [Wal18, Proposition 2.1].
Proposition 3.1.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a faithfully flat morphism. If $\operatorname{Ass}(IB)=\{{\mathfrak{p}}B~{}|~{}{\mathfrak{p}}\in\operatorname{Ass}(I)\}$, then $\rho(I)=\rho(IB)$.
Now we discuss some specific classes of morphisms which together with flatness gives faithfully flat. We start with the following definition.
Definition 3.2.
A ring homomorphism $\phi:A\to B$ has prime extension property if for every prime $\mathfrak{p}$ of $A$, $\phi({\mathfrak{p}})=\mathfrak{p}B$ is also prime in $B$. We also call $\phi$ is a prime extension.
Remark 3.3.
We recall some facts about the morphisms with prime extension properties.
  1. (1)
    In many literature, (for example, see, [HJ21]) prime extension has been defined as either ${\mathfrak{p}}B$ is a prime ideal or, the unit ideal. For us, prime extension means that prime ideals extends to a prime ideals.
  2. (2)
    Flat and prime extension implies faithfully flatness. Indeed, by [Mat89, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
Corollary 3.4.
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. If $I\subset A$ is a proper ideal and $\phi$ has prime extension properties then $\rho(I)=\rho(IB)$ and $\rho_{a}(I)=\rho_{a}(IB)$.
Proof.
Since $\phi$ is prime extension then by Remark 3.3, it is faithfully flat. Thus using Theorem 2.5 we get that
$I^{(m)}\subset I^{r}\text{ if and only if }(IB)^{(m)}\subset(IB)^{r},\text{ for all }m,r.$
Hence
$\rho(I)=\rho(IB),\text{ and }\rho_{a}(I)=\rho_{a}(IB).$
Flat and prime extension has geometric meanings. Before we start to talk about them, we recall the definition.
Definition 3.5.
Let $f:X\to Y$ be a morphism of schemes, and let $y\in Y$ be a point. Let $\kappa(y)$ be the residue field of y, i.e. $\kappa(y)=\mathcal{O}_{y}/\mathfrak{m}_{y}$, and let $\operatorname{Spec}\,\kappa(y)\rightarrow Y$ be the natural morphism. Then we define the fiber of the morphism $f$ over the point $y$ to be the scheme
$X_{y}=X\times_{Y}\hskip 1.42262ptSpec\,\kappa(y).$
Note that fiber $X_{y}$ is a scheme over $\kappa(y)$.
Let $f:\operatorname{Spec}B\to\operatorname{Spec}A$ be a morphism of schemes, induced by ring homomorphism $\phi:A\to B$. Then the fiber over $\mathfrak{p}\in\operatorname{Spec}A$ is given by $\operatorname{Spec}B_{\mathfrak{p}}=\operatorname{Spec}\left(B\otimes_{A}\kappa(\mathfrak{p})\right)=\operatorname{Spec}\left(B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}\right)$. We have the following proposition:
Proposition 3.6.
If the ring homomorphism $\phi:A\to B$ is flat then $B_{\mathfrak{p}}$ is an integral domain if only if $\mathfrak{p}B$ prime.
Proof.
Since $\phi$ is flat, we can consider $\mathfrak{p}\otimes_{A}B$ as $\mathfrak{p}B$. It’s easy to see the followings:
$B_{\mathfrak{p}}\text{ is an integral domain }\iff B/\mathfrak{p}B\text{ is anintegral domain }\iff\mathfrak{p}B\text{ is prime}.$
This proposition implies that if $A$ is reduced, and $\phi$ has prime extension property, then $B$ is also reduced. Indeed, if $A$ contains the zero ideal as a prime ideal, so does $B$. It also gives us an enormous class of examples. If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain, which is geometrically interesting. Before we give an example, we recall the following lemma.
Lemma 3.7.
[Sta24, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
Example 3.8.
Let ${\mathbbm{k}}$ be an algebraically closed field, and let
$A={\mathbbm{k}}[t]\text{ and }B={\mathbbm{k}}[t,x,y]/(y-x^{2}-tx).$
Let $\phi$ to be the ring homomorphism from $A$ to $B$ by sending $t$ to $t$, and let $\Phi$ be the morphism from $\operatorname{Spec}B$ to $\operatorname{Spec}A$ induced by $\phi$. Note that $A\otimes_{A}B$ is a domain. It is not hard to see $\phi$ is a flat morphism. The map from $I\otimes_{A}B$ to $A\otimes_{A}B$ is always injective for every ideal $I\subset A$, then by Lemma 3.7 we get that $\phi$ is flat. The injectivity comes from the fact that both $y-x^{2}$ and $y-x^{2}-tx$ are irreducible in ${\mathbbm{k}}[t,x,y]$.
We identify closed points of $A$ with elements of ${\mathbbm{k}}$, and obviously the $(t-c)$ in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension. The fiber $\operatorname{Spec}B_{(t-c)}$ is given by the plane curve $y=x^{2}+cx$ in $\mathbb{A}_{\mathbbm{k}}^{2}$, which is an irreducible variety, and those plane curves are even homeomorphic to each other.
The situation doesn’t always perform so nicely. If we consider $A$ as ${\mathbbm{k}}[t,s]$ and $B$ as ${\mathbbm{k}}[t,s,x,y]/(y-x^{2}-tx)$, we also have $(t-c)$ as our point, then $B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ will be ${\mathbbm{k}}[s,s^{-1},x,y]/(y-x^{2}-cx)={\mathbbm{k}}(s)[x,y]/(y-x^{2}-cx)$. However we have $(t-a,s-b)$ as our point where $a$,$b\in{\mathbbm{k}}$, now the corresponding fiber is $\operatorname{Spec}{\mathbbm{k}}[x,y]/(y-x^{2}-ax)$, which is not homeomorphic to the previous one.
4. Resurgence of sum of ideals
Reviewing the recent research of resurgence number of sums of ideals for polynomial rings over fields, our goal is to prove these results for sums of finitely generated ${\mathbbm{k}}$-algebras which are integral domains where ${\mathbbm{k}}$ is algebraically closed. In [BHJT21, Theorem 2.6] it has been proved that if $I$ and $J$ are ideals in different polynomial rings then
$\rho_{a}(I+J)=\overline{\rho}_{\alpha}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}.$
In [KNT24, Theorem 2.3], it has been proved that
$\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
In this section we prove similar results for finitely generated ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field. One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras. The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up. To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals.
Lemma 4.1.
Let $A$ and $B$ be finitely generated ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed, which are integral domains. If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes B$, then $\phi$ is flat and has prime extension property. Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$.
Proof.
Recall that the canonical ring homomorphism is defined by
$\phi:A\to A\otimes_{\mathbbm{k}}B\text{ defined by }a\rightarrow a\otimes_{\mathbbm{k}}1.$
Observe that if we can show $\phi$ is injective, then $I\rightarrow A\otimes_{A}R$ is injective for every ideal $I\subset A$. Hence by Lemma 3.7, $\phi$ is flat. Injectivity of $\phi$ is obvious since both $A$ and $B$ are fintely generated algebras over an algebraically closed field ${\mathbbm{k}}$, and integral domains. If there exists a nonzero element in ${\mathbbm{k}}$ annihilating some elements in $A$ (or in $B$), then this element has to be zero in $A$ (or in $B$), which is a contradiction.
Let $\mathfrak{p}\subset A$ be a prime ideal, we need to prove that ${\mathfrak{p}}R$ is also prime. Since $\phi$ is injective then $\mathfrak{p}R=\mathfrak{p}\otimes_{\mathbbm{k}}B$. Observe that $R/\mathfrak{p}R=\left(A/\mathfrak{p}\right)\otimes_{\mathbbm{k}}B$ because $R$ is an integral domain.
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Corollary 3.4. ∎
Remark 4.2.
To see why we need the conditions on $A$ and $B$, we will need some examples and theorems. Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain.
The next lemma also extends [BHJT21, Lemma 2.4]. The proof of [BHJT21, Lemma 2.4] depends on divisibility in the polynomial ring. While working with general domains, we treated them as vector spaces and used properties of basis.
Lemma 4.3.
Let $A$ and $B$ be finitely generated ${\mathbbm{k}}$-algebras, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ be the sum of ideal extensions of $I$ and $J$ in $R=A\otimes_{\mathbbm{k}}B$ then
$\rho_{a}^{\lim\sup}(I+J)\geqslant{\max}\left\{\rho_{a}^{{\lim\sup}}(I),\,\rho_{a}^{{\lim\sup}}(J)\right\}.$
Proof.
Take $\theta\in{\mathbb{Q}}$ such that $\theta>\rho_{a}^{\lim\sup}(I+J)$. Then there is a $s_{0}\in{\mathbb{N}}$ such that for all $t\geqslant s_{0},\theta>\rho(I+J,t)$, where $\rho(I+J,t)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subset I^{r},m\geqslant t,r\geqslant t\right\}$. Thus for any $h,r\in{\mathbb{N}}$ so that $h,r\geqslant t\geqslant s_{0}$ and $\dfrac{h}{r}\geqslant\theta$, one has $(I+J)^{r}\supset(I+J)^{(h)}\supset I^{(h)}$. Up to this point the proof idea is same as [BHJT21, Lemma 2.4]. Now we need to show that $I^{(h)}\subseteq I^{r}$.
Let $V$ be a ${\mathbbm{k}}$ basis of $I$, which extends to $V^{*}$ a ${\mathbbm{k}}$-basis of $A$ containing $1$. Let $W$ be a ${\mathbbm{k}}$-basis of $J$, which extends to $W^{*}$, a ${\mathbbm{k}}$-basis of $B$ containing $1$. Then $V\otimes W^{*}$ is a ${\mathbbm{k}}$-basis for $I\otimes B$ and $V^{*}\otimes W$ is a ${\mathbbm{k}}$-basis for $V\otimes W^{*}$. Also, $V^{*}\otimes W^{*}$ is a ${\mathbbm{k}}$-basis of $R=A\otimes B$. Let $f\in I^{(h)}$ then $f=\sum_{i=1}^{n}c_{i}v_{i}$, where $c_{i}\in{\mathbbm{k}}\setminus\{0\}$ and $v_{i}\in V$. Now
$f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(I+J)^{(h)}\subset(I+J)^{r}=\sum_{j=0}^{r}(IR)^{i}(JR)^{r-j}=\sum_{j=0}^{r}\left((IR)^{j}\cap(JR)^{r-j}\right).$
The last equality follows from [HNTT20, Lemma 3.1]. Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes_{k}1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$. Since $1\notin W$ then $v\otimes 1\notin A\otimes J$. Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $11\leqslant j\leqslant r-$ 1. Hence $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(IR)^{h}$. Now $(IR)^{h}\cong I^{h}\otimes B$ and thus $f\otimes 1\in(IR)^{h}$ implies that $f\in I^{h}$. This completes the proof. ∎
The following generalizes [BHJT21, Lemma 2.5] to finitely generated algebra domains. The proof of [BHJT21, Lemma 2.5] depends on the binomial summation formula and numerical inequalities, which holds in general domains as well, so we omit the proof.
Lemma 4.4.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ denotes $IR+JR$, where $R=A\otimes_{\mathbbm{k}}B$, then $\rho_{a}(I+J)\leqslant\max\left\{\rho_{a}^{\lim\sup}(I),\rho_{a}^{\lim\sup}(J)\right\}$.
Now we give the formula for the asymptotic resurgence of the sum of ideals in finitely generated ${\mathbbm{k}}$-algebra domians where ${\mathbbm{k}}$ is an algebraically closed field.
Theorem 4.5.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. If $I+J$ denotes $IR+JR$, where $R=A\otimes_{\mathbbm{k}}B$, then
$\rho_{a}(I+J)=\overline{\rho}_{\alpha}(I+J)=\overline{\rho}(I+J)={\max}\{\rho_{a}(I),\rho_{a}(J)\}=\max\{\rho_{a}(IR),\rho_{a}(JR)\}.$
Proof.
The proof has the same spirit as [BHJT21, Theorem 2.6] and it follows from Lemma 4.1, Lemma 4.3 and Lemma 4.4. ∎
Now we give the result regarding the summation formula for resurgence of ideals in finitely generated ideals in ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed field.
Theorem 4.6.
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. Set $I+J=IR+JR$, where $R=A\otimes B$. Then
$\max\{\rho(I),\,\rho(J)\}\leqslant\rho(I+J)\leqslant\max\left\{\rho(I),\,\rho(J),\,\frac{2(\rho(I)+\rho(J))}{3}\right\}.$
Furthermore, if $\max\{\rho(I),\,\rho(J)\}\geqslant 2\min\{\rho(I),\,\rho(J)\}$, then $\rho(I+J)=\max\{\rho(I),\,\rho(J)\}$.
Proof.
We give a sketch of the proof as it follows the same pattern as [KNT24, Theorem 2.3]. First note that by Lemma 4.1, $\rho(IR)=\rho(I)$ and $\rho(JR)=\rho(J)$. The binomial summation formula for symbolic powers hold for ideals in noetherian ${\mathbbm{k}}$-algebras (see [HJKN23, Theorem 4.1]. The other numerical inequalities in [KNT24, Theorem 2.3] follows straightforward. Hence the required result follows. ∎